OpenStudy (laddiusmaximus):

Can someone please explain this step by step? Find the derivative of x^(2) +3x - 5 using the definition of the derivative (i.e.) using limits. I know the answer is 2x+3 using power rules and the like, but I can't figure out how and where to plug in the numbers using f(x) (x+h)-f(x)/h

OpenStudy (anonymous):

\[f(x)=x^2+3x-5\] \[f(x+h)=(x+h)^2+3(x+h)-5=x^2+2xh+h^2+3x+3h-5\] \[f(x+h)-f(x)=2xh+h^2+3h\] \[\frac{f(x+h)-f(x)}{h}=2x+3+h^2\] now take the limit as h goes to zero (by replacing h by zero) and get \[2x+3\]

OpenStudy (laddiusmaximus):

ok why is it 2xh?